OpenStudy (anonymous):

Hi! I have a problem that I am having trouble with. It says to find a equation of the line containing the given pair of points. (-3,0) (0,4)

OpenStudy (anonymous):

Do you know how to find slope?

hero (hero):

\[\frac {y_2 - y_1} {x_2 - x_1}\]

OpenStudy (anonymous):

Yes I know how to find the slope

OpenStudy (anonymous):

What did you get for slope?

OpenStudy (anonymous):

3/4

hero (hero):

Hmmm, I got 4/3

OpenStudy (anonymous):

y- y1=m(x- x1) Of the two points that you have (-3,0), (0,4) choose one: then plug in y value of that point into y1 plug in x value of into x1 plug in slope into m

OpenStudy (anonymous):

I do not get it

hero (hero):

\[y - 0 = \frac {4} { 3}(x -(-3))\]\[y - y_1 = m(x - x_1) \]

OpenStudy (anonymous):

is y the slope or what?

hero (hero):

\[y = \frac {4}{3}x+\frac {13}{3} \]

hero (hero):

m = slope

OpenStudy (anonymous):

then how do I figure out the value of y?

hero (hero):

You don't.....you only enter the points for \[x_1, y_1\] and m and then simplify from there.

OpenStudy (anonymous):

y=mx+b?

OpenStudy (anonymous):

the answer is above me, isnt it?

OpenStudy (anonymous):

hero already answered

hero (hero):

Yes

OpenStudy (anonymous):

y=4/3x+13/3

OpenStudy (anonymous):

is that it hero?

hero (hero):

Yes